# Growth of Gravitational Wave Spectrum from Sound Waves in a Universe with Generic Expansion Rate

### v1

@@ Abstract
We derived here the factor $\Upsilon$, which quantifies how the gravitational wave spectrum generated by sound waves in the radiation sector grows over time, in a universe with a generic expanding rate set by another dominant energy content. When the dominant energy density satisfies $\rho\propto a^{-3(1+w)}$, we found that $\Upsilon$ has a compact analytical expression: $\Upsilon=\frac{2[1-y^{3(w-1)/2}]}{3(1-w)}$, where $y=a(t)/a(t_{s})$ which is the ratio of the scale factor at a later time $t$ to that at $t_{s}$ when gravitational wave production from sound waves starts. This generic result reduces to that derived previously for radiation-dominated and matter-dominated cases, thus generalizing previous formulas to more general cosmological contexts and providing more accurate results. The derivation relies solely on a stationary source, implying that this generic result of $\Upsilon$ serves as an universal factor in describing the growth of the gravitational wave production and can appear beyond cosmological phase transitions.
@@ I Introduction
The stochastic background of gravitational waves from cosmological first-order phase transitions is one of the promising frontiers for discovering physics beyond the Standard Model. LIGO data has been used to search for gravitational waves originating from new physics at energy scales of $\mathcal{O}(10^{3}\sim 10^{6})$ TeV Romero et al. (2021); Badger et al. (2023), far exceeding the energy scale accessible by the current large colliders. Additionally, pulsar timing array experiments reported evidences for a stochastic gravitational wave background, indicating QCD phase transitions or strongly supercooled phase transitions as possible sources Xue et al. (2021); Arzoumanian et al. (2021); Agazie et al. (2023); Antoniadis et al. (2023, 2024); Xiao et al. (2023); Han et al. (2024); Ahmadvand et al. (2023); Zheng et al. (2024); Bian et al. (2021, 2024). Space-based gravitational waves detectors, such as LISA Amaro-Seoane et al. (2017), Taiji Hu and Wu (2017); Ruan et al. (2020) and Tianqin Luo et al. (2016); Mei et al. (2021); Luo et al. (2020), are currently in preparation and expected to come into play in the coming decades.
@@
To ascertain whether the detected gravitational waves indeed originated from the early universe phase transitions, a profound understanding of the relevant processes is essential. It is widely believed that phase transition gravitational waves can primarily be generated through three mechanisms: bubble wall collisions Kosowsky and Turner (1993); Kosowsky et al. (1992); Jinno and Takimoto (2017); Zhao et al. (2023); Di et al. (2021); Li et al. (2023), sound waves Hindmarsh et al. (2014, 2015), and magnetohydrodynamic (MHD) turbulence Caprini et al. (2009); Kahniashvili et al. (2008, 2010); Caprini et al. (2016); Roper Pol et al. (2020). Gravitational waves produced by bubble walls can be well-described by the envelope approximation, and an analytical expression has been obtained Jinno and Takimoto (2017). In thermal plasma phase transitions, gravitational waves induced by sound waves are dominant and have been extensively studied, while those from MHD turbulence are sub-dominant and related studies remain highly uncertain. Although the behavior of gravitational waves induced by sound waves can be reasonably described by the sound shell model, simulation results suggest the presence of additional mechanisms affecting the final spectrum Cutting et al. (2020). This indicates a need for further investigation into this process. Recently, Ref. Cai et al. (2023) focused on the forced propagating contribution from the initial collision stage of sound shells and achieved better agreement with simulations compared to the sound shell model. Ref. Ellis et al. (2020) compared several models and discovered that the parameter regions with high signal-to-noise ratios (SNR) often coincide with short duration of sound waves, potentially reducing the observable gravitational wave spectrum. In our previous work Guo et al. (2021), we more carefully calculated the gravitational wave spectrum of the sound shell model, and was able to extract an overall factor connected with the duration of sound waves, called $\Upsilon$. Ref. Roper Pol et al. (2024) pointed out that some assumptions in the sound shell model lead to incorrect infrared behavior and extended the $\Upsilon$ factor to the infrared regime. However, most of the above calculations are in the radiation-dominated universe and have not been verified in other universe scenarios. If the equation of state of the universe deviates from the ultra-relativistic case during a phase transition, the gravitational wave spectrum may be modified Giombi et al. (2024). Therefore, it is both necessary and interesting to extend these results to a more general cosmological period to reproduce the potential gravitational wave generation processes as accurately as possible.
The aim of this work is to follow our previous work and try to extend the $\Upsilon$ factor from the radiation-dominated universe and the matter-dominated universe to an arbitrary single-component dominated universe. The structure of the paper is as follows: in section II, we present the theoretical framework for the sound shell model, followed by our extensions in section III. In section IV, we provide the resulting gravitational wave spectrum, and finally, the conclusions are discussed in section V.
II Theoretical Framework and Sound Shell Model
@@
We first consider the gravitational waves in the FLRW universe, described by the metric
@@
$\dot{h}_{ij}(t,\mathbf{q}_{1})\dot{h}_{ij}(t,\mathbf{q}_{2})\right>=(2\pi)^{2}\delta^{3}(\mathbf{q}_{1}+\mathbf{q}_{2})P_{\dot{h}}(q_{1},t).$ (4)
@@
The gravitational wave equation Eq.(2) could be solved by the Green’s function with the following boundary conditions
$G(\tilde{\eta}\leq\tilde{\eta}_{0})=0,$ (8)
$\left.\frac{\partial G(\tilde{\eta},\tilde{\eta}_{0})}{\partial\tilde{\eta}}\right|_{\tilde{\eta}=\tilde{\eta}_{0}^{+}}=1,$ (9)
where $\tilde{\eta}=q\eta$ is a dimensionless quantity and the subscript zero indicates the time when the phase transition begins. The solution is then given by
$h_{ij}(t,\mathbf{q})=16\pi G\int_{\tilde{\eta}_{0}}^{\tilde{\eta}}{\rm d}\tilde{\eta}^{\prime}G(\tilde{\eta},\tilde{\eta}^{\prime})\frac{a^{2}(\tilde{\eta}^{\prime})\pi_{ij}^{T}(\tilde{\eta}^{\prime},q)}{q^{2}}.$ (10)
Inserting the above solution into the left hand of Eq.(4), we could obtain
@@
×$\left.~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\times\frac{a^{2}(\eta_{1})a^{2}(\eta_{2})}{q^{2}}\Pi^{2}(q_{1},\eta_{1},\eta_{2})\right\},$ (11)
@@
$\pi_{ij}^{T}(\eta_{1},\mathbf{q}_{1})\pi_{ij}^{T}(\eta_{2},\mathbf{q}_{2})\right>=(2\pi)^{3}\delta^{3}(\mathbf{q}_{1}+\mathbf{q}_{2})\Pi^{2}(q_{1},\eta_{1},\eta_{2}).$ (12)
The transverse-traceless part of the energy-momentum tensor, $\pi_{ij}^{T}$, is given by
@@
where $\Lambda_{ij,kl}$ is the transverse-traceless projection operator, $\bar{p}$ and $\bar{e}$ are the averaged pressure and energy density. Then we could express the $\Pi^{2}$ as the four point correlation function of the fluid velocity. With the following Fourier transformation of the Gaussian random field of the fluid velocity $\tilde{v}^{i}_{\mathbf{q}}(\eta)$,
$\tilde{v}_{\mathbf{q}}^{i}(\eta)=\int{\rm d}^{3}x~{}v^{i}(\mathbf{x},\eta)e^{-i\mathbf{q}\cdot\mathbf{x}},\$ (14)
and according to the Wick theorem, $\Pi^{2}$ could be simplified as follows
$\Pi^{2}(k,\eta_{1}-\eta_{2})=\left(\bar{e}+\bar{p}\right)^{2}\int$ $\frac{{\rm d}^{3}q}{(2\pi)^{3}}\bigg{\{}G(q,\eta_{1},\eta_{2})G(\tilde{q},\eta_{1},\eta_{2})$
×$\left.\times\frac{q^{2}}{\tilde{q}^{2}}(1-\mu^{2})^{2}\right\},$ (15)
where $\tilde{q}=\left|\mathbf{q}-\mathbf{k}\right|$ and $\mu=\hat{\mathbf{q}}\cdot\hat{\mathbf{k}}$. In the derivation of the above expression, we have used the UETC of the Fourier coefficients of fluid velocity
$\tilde{v}^{i}_{\mathbf{q}}(\eta_{1})\tilde{v}^{j*}_{\mathbf{k}}(\eta_{2})\right>=\delta^{3}(\mathbf{q}-\mathbf{k})\hat{q}^{i}\hat{k}^{j}G(q,\eta_{1},\eta_{2}),$ (16)
and the property of the projection operator
$\Lambda_{ij,kl}\hat{q}^{i}\hat{\tilde{q}}^{j}\hat{q}^{k}\hat{\tilde{q}}^{l}=\frac{q^{2}}{\tilde{q}}\frac{(1-\mu^{2})^{2}}{2}.$ (17)
Note that the source of the gravitational wave here is approximately stationary. This implies that $\Pi^{2}$ only depends on $\eta_{1}-\eta_{2}$, and leads to a factorized form for the spectrum, corresponding to the $\Upsilon$ factor Guo et al. (2021).
@@
In a first-order phase transition, the expanding vacuum bubble wall exerts an influence on the surrounding fluid. After bubble collisions, we expect the fluid to propagate like sound waves and become the dominant source for gravitational waves production. The calculation of $G(q,\eta_{1},\eta_{2})$ requires obtaining the specific form of $\tilde{v}_{\mathbf{q}}^{i}$ after bubble collision, in other words, we need to establish a relationship between the fluid velocity before the bubble collision ($\eta<\eta_{c}$) and that after the bubble collision ($\eta>\eta_{c}$). Because the vacuum bubbles are nucleated randomly, it is easier to look at only one bubble first, for example the $n$-th bubble. We could formally decompose this velocity field before collision as
$v^{i(n)}(\eta\leq\eta_{c}^{(n)},\mathbf{x})$ $=\frac{1}{2}\int\frac{{\rm d}^{3}q}{(2\pi)^{3}}\left[\tilde{v}^{i(n)}_{\mathbf{q}}(\eta)e^{i\mathbf{q}\cdot\mathbf{x}}\right.$
$\left.~{}~{}~{}~{}~{}~{}~{}~{}+\tilde{v}^{i(n)*}_{\mathbf{q}}(\eta)e^{-i\mathbf{q}\cdot\mathbf{x}}\right].$ (19)
After bubble collision, the sound waves freely propagate and admit the following decomposition
$v^{i(n)}(\eta\geq\eta_{c}^{(n)},\mathbf{x})$ $=\int\frac{{\rm d}^{3}q}{(2\pi)^{3}}\left[v_{\mathbf{q}}^{i(n)}e^{-i\omega\eta+i\mathbf{q}\cdot\mathbf{x}}\right.$
$\left.~{}~{}~{}~{}~{}~{}~{}~{}+v_{\mathbf{q}}^{i(n)*}e^{i\omega\eta-i\mathbf{q}\cdot\mathbf{x}}\right],$ (20)
where $\omega=c_{s}q$. To match the above two different processes, it is natural to evaluate those equations at collision time $\eta_{c}$ simultaneously and we will find that
$v_{\mathbf{q}}^{i(n)}=\frac{1}{2}\left[\tilde{v}_{\mathbf{q}}^{i(n)}(\eta_{c}^{(n)})+\frac{i}{\omega}\tilde{v}_{\mathbf{q}}^{i(n)\prime}(\eta_{c}^{(n)})\right]e^{i\omega\eta_{c}^{(n)}}.$ (21)
By the Euler equation of the fluid
$\tilde{v}_{\mathbf{q}}^{j\prime}+ic_{s}^{2}q^{j}\tilde{\lambda}_{\mathbf{q}}=0,$ (22)
where $\tilde{\lambda}_{\mathbf{q}}$ is the Fourier coefficients of the energy fluctuation $\lambda=\left[\tilde{e}(x)-\bar{\tilde{e}}\right]/\bar{\tilde{w}}$, we could write the Fourier coefficients of the sound waves as
$v_{\mathbf{q}}^{i(n)}=\frac{1}{2}\left[\tilde{v}_{\mathbf{q}}^{i(n)}(\eta_{c}^{(n)})+c_{s}\hat{q}^{i(n)}\tilde{\lambda}_{\mathbf{q}}(\eta_{c}^{(n)})\right]e^{i\omega\eta_{c}^{(n)}}.$ (23)
According to the assumption of the sound shell model, the total velocity and energy density field are the superposition of the corresponding fields produced by the $N_{b}$ expanding bubblesHindmarsh (2018); Hindmarsh and Hijazi (2019)
$v_{\mathbf{q}}^{i}=\sum_{n}^{N_{b}}v_{\mathbf{q}}^{i(n)}.$ (24)
The velocity field configuration produced by a single bubble before bubble collision can be obtained using the standard energy budget method Espinosa et al. (2010). It is straightforward to relate this to its corresponding Fourier coefficients
$\tilde{v}_{\mathbf{q}}^{j(n)}(\eta_{c}^{(n)})$ $=ie^{-i\mathbf{q}\cdot\mathbf{x}^{(n)}}\left[T^{(n)}\right]^{3}\hat{z}^{(n)j}f^{\prime}(z^{(n)}),$ (25)
$\tilde{\lambda}_{\mathbf{q}}^{(n)}(\eta_{c}^{(n)})$ $=e^{-i\mathbf{q}\cdot\mathbf{x}^{(n)}}\left[T^{(n)}\right]^{3}l(z^{(n)}),$ (26)
where $T^{(n)}\equiv\eta_{c}^{(n)}-\eta_{\rm nucleate}^{(n)}$ represents the lifetime of the $n$-th bubble, and $z^{n}\equiv\mathbf{q}T^{(n)}$. The specific form of $f(z)$ and $l(z)$ can be seen in Ref.Guo et al. (2021); Hindmarsh and Hijazi (2019). Inserting those into Eq.(23), we finally get
$v_{\mathbf{q}}^{i(n)}=i\hat{z}^{(n)i}\left[T^{(n)}\right]^{3}e^{i\omega\eta_{c}^{(n)}-i\mathbf{q}\cdot\mathbf{x}^{(n)}}A(z^{n}),$ (27)
where $A(z)=\left[f^{\prime(z)}-ic_{s}l(z)\right]/2$.
Once we have obtained $v^{i(n)}_{\mathbf{q}}$, we could directly compute $\tilde{v}_{\mathbf{q}}^{i}(\eta\geq\eta_{c}^{(n)})$ by its definition. After doing the ensemble average over nucleation sites of all bubbles, the velocity power spectrum $G(q,\eta_{1},\eta_{2})$ is
$G(q,\eta_{1},\eta_{2})=2P_{v}(q){\rm cos}\left[\omega(\eta_{1}-\eta_{2})\right],$ (28)
where $P_{v}$ contains the integration of $A(z)$, and its specific form can be also found in Ref.Guo et al. (2021); Hindmarsh and Hijazi (2019). It is worth noting that Ref.Roper Pol et al. (2024) pointed out that there is an extra term in Eq.(28), which is important to reproduce the causality condition in the infrared region.
III $\Upsilon$ factor in singlet component universe
Now we have gathered all the components except for the Green’s function. Since we aim to obtain the $\Upsilon$ factor for arbitrary single-component background universes, we need to extend the calculation of the Green’s function into a more general case.
By the continuity equation for a perfect fluid
@@ III $\Upsilon$ factor in singlet component universe
Inserting Eq.(30) into the Friedmann equation of flat universe
$\frac{{\rm d}a}{{\rm d}\eta}=a^{2}\sqrt{\frac{8\pi G}{3}\rho}=a^{2}\sqrt{H_{s}^{2}\left(\frac{a}{a_{s}}\right)^{-3(1+w)}},$ (31)
and defining $y(\eta)=a(\eta)/a_{s}$, following our previous derivation Guo et al. (2021), we could obtain
$\frac{{\rm d}y}{{\rm d}\eta}=a_{s}H_{s}y^{(1-3w)/2},$ (32)
where $H_{s}$ is the Hubble constant when the source begins active. With the initial condition $y(\eta_{s})=1$, the solution of the above equation is
$y^{\frac{1+3w}{2}}=\frac{3w+1}{2}(\eta-\eta_{s})H_{s}a_{s}+1.$ (33)
Changing the variable from $\eta$ to $y$, the gravitational wave equation Eq.(2) could be rewritten as
@@
by setting $2n=2-n(3w-1)$. The corresponding Green’s function is the homogeneous solution of this equation and could be expressed as Bowman (2012)
@@
where $J_{\nu}(x)$ is the Bessel function of the first kind and the $\nu=(3w-3)/(6w+2)$ should not be an integer. If $\nu$ is an integer, the $J_{-\nu}$ in the above expression should be replaced with the Bessel function of the second kind to satisfy the condition of linear independence Bowman (2012). In this paper, we confine ourselves to values of $w$ that satisfy the strong-energy condition, i.e., $01+3w>$ 0, which implies that almost all values of $\nu$ are not integers. The constants $C_{1}$ and $C_{2}$ can be determined by the following initial conditions
$G(\tilde{y}\leq\tilde{y}_{0})=0,$ (37)
$\left.\frac{\partial G(\tilde{y},\tilde{y}_{0})}{\partial\tilde{y}}\right|_{\tilde{y}=\tilde{y}_{0^{+}}}=\tilde{y}_{0}^{3w+1},$ (38)
with the following results:
$C_{1}$ $=-\frac{b}{b_{1}a-ba_{1}}\tilde{y}_{0}^{\frac{3w+1}{4}},$ (39)
$C_{2}$ $=\frac{a}{b_{1}a-ba_{1}}\tilde{y}_{0}^{\frac{3w+1}{4}},$ (40)
where $a=J_{\nu}\left(\frac{2}{3w+1}\tilde{y}_{0}^{\frac{3w+1}{2}}\right)$, $b=J_{-\nu}\left(\frac{2}{3w+1}\tilde{y}_{0}^{\frac{3w+1}{2}}\right)$ and $a_{1}$ ($b_{1}$) is the corresponding derivative with respect to its variable, ${\rm d}J_{\nu}(z)/{\rm d}z$ (${\rm d}J_{-\nu}(z)/{\rm d}z$).
@@
×$\times\int_{y_{s}}^{y}{\rm d}y_{1}\int_{y_{s}}^{y}{\rm d}y_{2}\left\{\left(\frac{q}{a_{s}H_{s}}\right)^{4n-2}\frac{\partial G\left(\tilde{y},\tilde{y}_{1}\right)}{\partial\tilde{y}}\frac{\partial G\left(\tilde{y},\tilde{y}_{2}\right)}{\partial\tilde{y}}\right.$
×$\left.~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\times\frac{1}{y_{1}^{2}y_{2}^{2}}\frac{\tilde{\Pi}^{2}\left(qR_{*},q|\eta_{1}-\eta_{2}|\right)}{\tilde{q}^{2}}\right\}.$ (41)
@@
The corresponding Jacobian determinant is $(y_{1}y_{2})^{\frac{1-3w}{2}}$ and the new limits of integration are
$0\leq y_{-}\leq(y^{\frac{3w+1}{2}}-1)\frac{2}{3w+1},$ (45)
$\frac{2}{3w+1}+\frac{y_{-}}{2}\leq y_{+}\leq-\frac{y_{-}}{2}+\frac{2}{3w+1}y^{\frac{3w+1}{2}},$ (46)
and
$(1-y^{\frac{3w+1}{2}})\frac{2}{3w+1}\leq y_{-}\leq 0,$ (47)
$-\frac{y_{-}}{2}+\frac{2}{3w+1}\leq y_{+}\leq\frac{y_{-}}{2}+y^{\frac{3w+1}{2}}\frac{2}{3w+1}.$ (48)
To explicitly show the $a^{-4}$ behavior of the gravitational radiation, we also denote
@@
Combining the above results, we could further simplify Eq.(III) as
@@
×$~{}~{}~{}~{}\times\int{\rm d}y_{-}\frac{\tilde{\Pi}^{2}\left(q|\eta_{1}-\eta_{2}|\right)}{\tilde{q}^{2}}\int{\rm d}y_{+}\frac{\mathcal{G}\left(\tilde{y},\tilde{y}_{1},\tilde{y}_{2}\right)}{y_{1}^{\frac{3+3w}{2}}y_{2}^{\frac{3+3w}{2}}}.$ (50)
@@
To calculate this complex integration, it is worth noting that the variable of the Green’s function Eq.(III) is the combination $\tilde{q}=q/H_{s}a_{s}=qR_{*}/H_{s}a_{s}R_{*}\sim\mathcal{O}(10^{3})qR_{*}$. This indicates that we could use the asymptotic expansion of the Bessel functions

### v2

@@ Abstract
We derive a compact analytical expression for the growth factor $\Upsilon$, which characterizes how the gravitational wave spectrum sourced by sound waves evolves in a universe with a generic expansion history. Assuming the dominant energy density scales as $\rho\propto a^{-3(1+w)}$, we obtain $\Upsilon=\frac{2[1-y^{3(w-1)/2}]}{3(1-w)}$, where $y=a(t)/a(t_{s})$ is the ratio of the scale factor at a later time $t$ to that at $t_{s}$ when gravitational wave production from sound waves starts. This general result reduces to known forms in radiation- and matter-dominated eras, thereby extending previous formulas to a broader class of cosmological backgrounds. The derivation assumes only that the source is stationary, making $\Upsilon$ a universal factor that captures gravitational wave growth in various scenarios—not limited to phase transitions.
@@
Introduction–The stochastic background of gravitational waves from cosmological first-order phase transitions is one of the promising frontiers for discovering physics beyond the Standard Model Athron et al. (2024); Roshan and White (2024). LIGO data has been used to search for gravitational waves originating from new physics at energy scales of $\mathcal{O}(10^{3}\sim 10^{6})$ TeV Romero et al. (2021); Badger et al. (2023), far exceeding the energy scale accessible by the current large colliders. Additionally, pulsar timing array experiments reported evidences for a stochastic gravitational wave background, indicating QCD phase transitions or strongly supercooled phase transitions as possible sources Xue et al. (2021); Arzoumanian et al. (2021); Agazie et al. (2023); Antoniadis et al. (2023, 2024); Xiao et al. (2023); Han et al. (2024); Ahmadvand et al. (2023); Zheng et al. (2024); Bian et al. (2021, 2024). Space-based gravitational waves detectors, such as LISA Amaro-Seoane et al. (2017), Taiji Hu and Wu (2017); Ruan et al. (2020) and Tianqin Luo et al. (2016); Mei et al. (2021); Luo et al. (2020), are currently in preparation and expected to come into play in the coming decades.
@@
To ascertain whether the detected gravitational waves indeed originated from the early universe phase transitions, a profound understanding of the relevant processes is essential. It is widely believed that phase transition gravitational waves can primarily be generated through three mechanisms: bubble wall collisions Kosowsky and Turner (1993); Kosowsky et al. (1992); Jinno and Takimoto (2017); Zhao et al. (2023); Di et al. (2021); Li et al. (2023), sound waves Hindmarsh et al. (2014, 2015), and magnetohydrodynamic (MHD) turbulence Caprini et al. (2009); Kahniashvili et al. (2008, 2010); Caprini et al. (2016); Roper Pol et al. (2020). Gravitational waves produced by bubble walls can be well-described by the envelope approximation, and an analytical expression has been obtained Jinno and Takimoto (2017). In thermal plasma phase transitions, gravitational waves induced by sound waves are dominant and have been extensively studied, while those from MHD turbulence are sub-dominant and related studies remain highly uncertain. Although the behavior of gravitational waves induced by sound waves can be reasonably described by the sound shell model, simulation results suggest the presence of additional mechanisms affecting the final spectrum Cutting et al. (2020). This indicates a need for further investigation into this process. Recently, Ref. Cai et al. (2023) focused on the forced propagating contribution from the initial collision stage of sound shells and achieved better agreement with simulations compared to the sound shell model. Ref. Ellis et al. (2020) compared several models and discovered that the parameter regions with high signal-to-noise ratios (SNR) often coincide with short duration of sound waves, potentially reducing the observable gravitational wave spectrum. In our previous work Guo et al. (2021), we calculated the gravitational wave spectrum of the sound shell model more carefully, and was able to extract an overall factor connected with the duration of sound waves, called $\Upsilon$. Ref. Roper Pol et al. (2024) pointed out that some assumptions in the sound shell model lead to incorrect infrared behavior and extended the $\Upsilon$ factor to the infrared regime. However, most of the above calculations are in the radiation-dominated universe and have not been verified in other universe scenarios. If the equation of state of the universe deviates from the ultra-relativistic case during a phase transition, the gravitational wave spectrum may be modified Giombi et al. (2024). Therefore, it is both necessary and interesting to extend these results to a more general cosmological period to reproduce the potential gravitational wave generation processes as accurately as possible.
The aim of this Letter is to follow our previous work and try to extend the $\Upsilon$ factor from the radiation-dominated universe and the matter-dominated universe to an arbitrary single-component dominated universe. We outline the main assumptions and derivation below, with technical details provided in the Supplemental Appendix.
@@
Conventions-We first consider the gravitational waves in the FLRW universe, described by the metric
@@
$\left<\dot{h}_{ij}(t,\mathbf{q}_{1})\dot{h}_{ij}(t,\mathbf{q}_{2})\right>=(2\pi)^{3}\delta^{3}(\mathbf{q}_{1}+\mathbf{q}_{2})P_{\dot{h}}(q_{1},t).$ (4)
@@
The gravitational wave equation Eq. (2) can be solved using the Green’s function method, allowing the spectrum function $P_{h^{\prime}}$ to be expressed as
@@
$\left.~~~~~~~~~~~~~~~~\times\frac{a^{2}(\eta_{1})a^{2}(\eta_{2})}{q^{2}}\Pi^{2}(q_{1},\eta_{1},\eta_{2})\right\},$ (8)
@@
$\left<\pi_{ij}^{T}(\eta_{1},\mathbf{q}_{1})\pi_{ij}^{T}(\eta_{2},\mathbf{q}_{2})\right>=(2\pi)^{3}\delta^{3}(\mathbf{q}_{1}+\mathbf{q}_{2})\Pi^{2}(q_{1},\eta_{1},\eta_{2}),$ (9)
$G$ is the Green’s function and $\tilde{\eta}=q\eta$ is dimensionless conformal time. For sound waves, the transverse-traceless part of the energy-momentum tensor, $\pi_{ij}^{T}$, is given by
@@
where $\Lambda_{ij,kl}$ is the transverse-traceless projection operator, $\bar{p}$ and $\bar{e}$ are the averaged pressure and energy density. In first-order phase transition, the fluid velocity $v_{i}$ in the vicinity of an individual vacuum bubble can be obtained from hydrodynamic. The full velocity field in space can then be modeled as a linear superposition of these individual velocity profiles. Under this case, correlation function $\Pi$ becomes approximately stationary, depending only on the time difference $\eta_{1}-\eta_{2}$. This leads to a factorized form of the gravitational wave spectrum
@@
$\Upsilon$ factor in singlet component universe-By the continuity equation for a perfect fluid
@@
Inserting Eq. (13) into the Friedmann equation for a flat universe and introducing a rescaled time variable $y(\eta)\equiv a(\eta)/a_{s}$, where the subscript $s$ denotes the moment when the source becomes active, Eq. (2) can be recast in the following form
@@
by setting 2n=2n(3w1)2n=2-n(3w-1). The corresponding Green’s function is the homogeneous solution of this equation and could be expressed as Bowman (2012)
@@
where $J_{\nu}(x)$ is the Bessel function of the first kind and the $\nu=(3w-3)/(6w+2)$ should not be an integer. If $\nu$ is an integer, the $J_{-\nu}$ in the above expression should be replaced with the Bessel function of the second kind to satisfy the condition of linear independence Bowman (2012). In this paper, we confine ourselves to values of $w$ that satisfy the strong-energy condition, i.e., $1+3w>0$, which implies that almost all values of $\nu$ are not integers.
@@
$\times\int_{y_{s}}^{y}{\rm d}y_{1}\int_{y_{s}}^{y}{\rm d}y_{2}\left\{\left(\frac{q}{a_{s}H_{s}}\right)^{4n-2}\frac{\partial G\left(\tilde{y},\tilde{y}_{1}\right)}{\partial\tilde{y}}\frac{\partial G\left(\tilde{y},\tilde{y}_{2}\right)}{\partial\tilde{y}}\right$.
$\left.~~~~~~~~~~~~~~~~~~~~~~~~\times\frac{1}{y_{1}^{2}y_{2}^{2}}\frac{\tilde{\Pi}^{2}\left(qR_{*},q|\eta_{1}-\eta_{2}|\right)}{\tilde{q}^{2}}\right\}.$ (17)
@@
In additional, to explicitly show the $a^{-4}$ behavior of the gravitational radiation, we also denote
@@
As a result, we could further simplify Eq.(Growth of Gravitational Wave Spectrum from Sound Waves in a Universe with Generic Expansion Rate) as
@@
$~~~~\times\int{\rm d}y_{-}\frac{\tilde{\Pi}^{2}\left(q|\eta_{1}-\eta_{2}|\right)}{\tilde{q}^{2}}\int{\rm d}y_{+}\frac{\mathcal{G}\left(\tilde{y},\tilde{y}_{1},\tilde{y}_{2}\right)}{y_{1}^{\frac{3+3w}{2}}y_{2}^{\frac{3+3w}{2}}}.$ (22)
@@
To calculate this complex integration, it is worth noting that the variable of the Green’s function Eq.(Growth of Gravitational Wave Spectrum from Sound Waves in a Universe with Generic Expansion Rate) is the combination $\tilde{q}=q/H_{s}a_{s}=qR_{*}/H_{s}a_{s}R_{*}\sim\mathcal{O}(10^{3})qR_{*}$. This indicates that we could use the asymptotic expansion of the Bessel functions